OpenStudy (anonymous):

Which expression is the simplified form of sqrt63W^36 ?

OpenStudy (anonymous):

\[\sqrt{63w^{36}}\]

OpenStudy (anonymous):

yea dat

OpenStudy (anonymous):

alright first take the number 63 and factor it to find what you can split it into. From doing so, you'll see that 63=9*7, correct?

OpenStudy (anonymous):

uh Yea

OpenStudy (anonymous):

next split the \[w^{36}\] into 2 terms. using \[w^n*w^d=w^{n+d}\] if you split it into two equal parts you get \[w^{18}*w^{18}\]

OpenStudy (anonymous):

now rewrite the square root \[\sqrt{7*9*w^{18}*w^{18}}\] the whole point of this is to try and get 2 terms inside the square root so you can pull them out.. another thing you can do is write 9 as 3*3 \[\sqrt{7*3*3*w^{18}*w^{18}}\]

OpenStudy (anonymous):

everything you have 2 of you pull out of the square root what do you get?

OpenStudy (anonymous):

so its 3W^18sqrt7?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[3w^{18}\sqrt{7}\]

OpenStudy (anonymous):

Thank You <3 !